OpenStudy (anonymous):

32-162k^2=0 solve the equation

OpenStudy (anonymous):

Why don't you show some work and let us check it for you?

OpenStudy (anonymous):

im checking my answers

OpenStudy (anonymous):

so what did you get?

OpenStudy (anonymous):

k= -1/8 1/8

OpenStudy (anonymous):

I think something went a little haywire... \[32-162k ^{2}=0\] add 162k^2 to both sides: \[32=162k ^{2}\] We can divide by 2: \[16=81k ^{2}\] Then take the square root of both sides...

OpenStudy (anonymous):

what you mean

OpenStudy (anonymous):

Do you understand what I did so far?

OpenStudy (anonymous):

4 and 9???

OpenStudy (anonymous):

Yes...4=+/-9k so then to solve for k you divide by 9, to get k=+/- 4/9